OpenStudy (anonymous):

Ok can someone give me the anwser to this so i will understand?! :) Solve the equation for x: 3x2 = 6x

OpenStudy (anonymous):

\[3x^2-6x=0\] is a start

OpenStudy (anonymous):

now factor as \[3x(x-2)=0\]

OpenStudy (anonymous):

set each factor equal to zero and solve \[3x=0\implies x=0\] \[x-2=0\implies x=2\] clear or need more explanation?

OpenStudy (anonymous):

i think i get it. so the anwser would be 2?

OpenStudy (anonymous):

and 0

OpenStudy (anonymous):

yes there are two answers since your original equation started with an \[x^2\]

OpenStudy (anonymous):

and you can check them both you started with \[3x^2=6x\] if you replace x by 0 you get \[0=0\] so you know that one works

OpenStudy (anonymous):

then if you replace x by 2 you get \[3\times 2^2=6\times 2\] \[12=12\] so that one works as well